Two independent random samples resulted in the following. Find the estimate for the standard error for the difference between two means. (Give your answer correct to two decimal places.) Sample A: nA = 21, sA = 8.6 Sample B: nB = 27, sB = 11.6
Answer #1

Solution,

Given that,

nA = 21

sA = 8.6

nB = 27

sB = 11.6

so,

SE ( xA, xB ) = (sA / nA ) + (sB / nB)

SE ( xA, xB ) = (8.6 / 21) + (11.6 / 27)

SE ( xA, xB ) = 0.40952381 + 0.42962963

SE ( xA, xB ) = 0.92